# Do LLMs have Consistent Values?

### v1

@@ Appendix E The Complete Item List of Best-Worst Refined Values (BWVr)
In our value anchoring approach, we used the description of values in [36] to prompt the LLMs. The set of descriptions is provided below.
@@ Appendix F Additional Results for Value Rankings
Table 3 provides additional results on value rankings for different prompting approaches.
Appendix G Cronbach $\alpha$ scores for temperature $0.7$
In the main text we provided results for Cronbach $\alpha$ with zero temperature. Here we provide further results for temperature $0.7$ in Figure 6. As expected, the numbers are overall lower than for the zero temperature case.
Figure 6: Comparing mean Cronbach’s $\alpha$ scores and standard deviations of LLMs in temperature 07 across different categories with human data [11].
Appendix H MDS figures for Gemini Pro and GPT 4 for temperature $0.0$
In the main text we provided the MDS plots for Gemini Pro for Value Anchor and Names. Here we provide further plots for Gemini Pro in Figure 7, and all the GPT 4 plots for temperature $0.0$ in Figure 8.
@@
Figure 7: Comparison of the MDS results between human data [11] and Gemini Pro for Demographic and Generated Persona respectively, in temperature $0.0$ condition.
@@
Figure 8: Comparison of the MDS results between human data [11] and GPT 4 for all prompts, in temperature $0.0$ condition.

### v3

@@ Appendix E The Complete Item List of Best-Worst Refined Values (BWVr)
In our value anchoring approach, we used the description of values in Lee et al. (2019) to prompt the LLMs. The set of descriptions is provided below.
@@ Appendix F Additional Results for Value Rankings
Table 2 provides additional results on value rankings for several prompting approaches.
Appendix G Additional MDS Plots
In the main text we provided the MDS plots for Gemini Pro for Value Anchor and Names. Here we provide further plots for Gemini Pro in Figure 7, all the GPT 4 plots in Figure 8, all of the Llama 3.1 8B plots in Figure 9 and all of Llama 3.1 70B plots Figure 10, and all of Gemma 2 9B plots Figure 11 and Gemma 2 27B plots Figure 12 for temperature $0.0$.
Appendix H Comparing Batch and Sequential Prompting
In the main text, we focused exclusively on batch prompting, where all items from the questionnaire were presented in a single prompt. An alternative is to present the questions in sequence, and ask the model to answer a question as soon as it is presented. To investigate potential differences between batch and sequential prompting, we evaluated on Llama models (in commercial models, sequential prompting is more expensive than batch). The value-ranking results are summarized in Figure 6, and the value-correlation results in Table 5. Regarding the values rankings, significant differences were observed between two datasets in some instances (e.g., Llama 3.1 70B, Basic: z = -2.64, p $=$ .004 ), while non-significant results were noted in other cases (e.g., Llama 3.1 8B, Value Anchor: z = -0.83, p $=$ .203). This indicates that, overall, the ranking correlations are closely aligned, with no clear inclination toward either batch or sequential prompting as better replicating the human value hierarchy. As for the value-correlations, it can be seen that the sequential prompts replicate the finding that the Value Anchor prompt best captures the circular structure of human values. Interestingly, for Llama 3.1 8B, batch prompting appeared to yield superior results. However, for Llama 3.1 70B, this was not the case across most prompts, suggesting that batch prompting may not consistently perform better across different models.
Figure 6: A heatmap of Spearman rank correlation between benchmark value hierarchies and dataset rankings for Llama 3.1 8B and 70B instruct for batch versus serial prompting methods, across temperature conditions.
Llama 3.1 8B
Value Anchor Demographic Generated Persona Names
Batch prompting
00 0.18 0.47 0.58 0.60
07 0.16 0.47 0.58 0.57
Serial prompting
00 0.18 0.54 0.65 0.61
07 0.18 0.54 0.65 0.37
Llama 3.1 70B
Value Anchor Demographic Generated Persona Names
Batch prompting
00 0.10 0.29 0.37 0.45
07 0.10 0.22 0.40 0.44
Serial prompting
00 0.14 0.26 0.20 0.48
07 0.14 0.26 0.20 0.69
@@ Appendix H Comparing Batch and Sequential Prompting
Figure 7: Comparison of the MDS results between human data (Schwartz and Cieciuch, 2022) and Gemini Pro for Demographic and Generated Persona respectively, in the temperature $0.0$ condition.
@@
Figure 8: Comparison of the MDS results between human data (Schwartz and Cieciuch, 2022) and GPT 4 for all prompts, in the temperature $0.0$ condition.
(a) Value Anchor
(b) Names
(c) Demographic
(d) Generated Persona
Figure 9: Comparison of the MDS results between human data (Schwartz and Cieciuch, 2022) and Llama 3.1 8B for all prompts, in the temperature $0.0$ condition.
(a) Value Anchor
(b) Names
(c) Demographic
(d) Generated Persona
Figure 10: Comparison of the MDS results between human data (Schwartz and Cieciuch, 2022) and Llama 3.1 70B for all prompts, in the temperature $0.0$ condition.
(a) Value Anchor
(b) Names
(c) Demographic
(d) Generated Persona
Figure 11: Comparison of the MDS results between human data (Schwartz and Cieciuch, 2022) and Gemma 2 9B for all prompts, in the temperature $0.0$ condition.
(a) Value Anchor
(b) Names
(c) Generated Persona
Figure 12: Comparison of the MDS results between human data (Schwartz and Cieciuch, 2022) and Gemma 2 27B for all prompts, in the temperature $0.0$ condition, with the exception of the Demographic prompt-(see Footnote 2).